In these two problems, assume all machinery was purchased at the first of the year.

                                                                                                                                                                        

Machine                                  Cost                 Salvage Value              Useful Life

1) Tractor A                             50,000               20,000                                   5

2) Tractor B                             60,000                  0                                         5

3) Combine                             84,000                   0                                         7

________________________________________________________________

  1. Assume you own the machinery above.  Calculate annual depreciation using the straight-line method.

(10 points)

Tractor A

Tractor B

Combine

Year 1

Year 2

Year 3

Year 4

Year 5

Year 6

Year 7

            

  1. Now calculate annual depreciation on the same equipment using double declining balance. Be careful not to exceed the salvage value.  If the salvage value is zero, switch to straight-line in the year when straight-line yields higher depreciation.  (Use the remaining value as the starting point when you change.)

(10 points)

Tractor A

Tractor B

Combine

Year 1

Year 2

Year 3

Year 4

Year 5

Year 6

Year 7
Answer #1

1. Straight line depreciation charge the same depreciation every year.

Per year depreciation = ( cost - salvage value ) / useful life.

For Tractor A : Depreciation every year = ( 50,000 - 20,000 ) /5 = 30000/5 = $ 6000 every year

For Tractor B : Depreciation every year = ( 60,000 -0 ) / 5. = $ 60,000/5 = $ 12000 every year.

In same way for combine : Depreciation every year = ( 84000 -0 ) /7. = $ 84,000 /7 = $ 12,000.

Thus,.

Tractor A Tractor B Combine
Year 1 $ 6000 $ 12,000 $ 12,000
Year 2 $ 6000 $ 12,000 $ 12,000
Year 3 $ 6000 $ 12,000 $ 12000
Year 4 $ 6000 $ 12,000 $ 12,000
Year 5 $ 6000 $ 12,000 $ 12000
Year 6 $ 12000
Year 7 $ 12000

2.

Double decline rate is double of straight line, and depreciation every year is charged on the latest book value at the beginning of year after adjusting accumulated depreciation from cost every year.

For Tractor A ,

Depreciation rate = 2 × ( 1/5) = 2 × 20% = 40%.

Depreciation for year 1 = Cost × rate = $ 50000 × 40 % = $ 20,000.

Depreciation for year 2 = revised cost × 20% = ( 50000-20000) × 20% = 30,000 × 40 % = $ 12,000.

But, the depreciation charge can not be more than the salvage value, so year 2 depreciation will be restricted to $ 10000 only , so that will lead to accumulated depreciation at $ 30,000 And cost will down to salvage value.

For Tractor B,

Depreciation rate = 2 × ( 1/5) = 40%.

Depreciation for year 1 = Cost × Rate = $ 60,000 × 40% = $ 24,000.

Depreciation for year 2 = revised cost × rate = ( 60,000 - 24,000) × 40% = $ 36000 × 40% = $ 14,400

Accordingly the depreciation for rest of the years can be calculated.as year 3 = $ 8640 , year 4 = $ 5184 and for year 5 = $ 3110.

For Combine:

Rate = 2 × (1/7) = 28.57%

Depreciation for year 1 = Cost × rate = $ 84,000 × 28.57% = $ 24,000.

Thus revised cost for year 2 depreciation = $ 84,000 - $ 24,000 = $ 60,000.

Year 2 depreciation = $ 60,000 × 28.57% = $ 17,143.

Accordingly the rest of years depreciation can be calculated which is shown in the table below.

Tractor A Tractor B Combine
Year 1 $ 20,000 $ 24,000 $ 24,000
Year 2 $ 10,000 $ 14,400 $ 17143
Year 3 $ 8640 $ 12245
Year 4 $ 5184 $ 8746
Year 5 $ 3110 $ 6247
Year 6 $ 4462
Year 7 $ 3187